Ask your own question, for FREE!
Mathematics 12 Online
OpenStudy (anonymous):

There are 12 girls and 14 boys in math class. The teacher puts the names of the students in a hat and randomly picks one name. Then the teacher picks another name without replacing the first. What is the probability that both students picked are boys?

OpenStudy (xguardians):

12 + 14 = 26 Students in Total. 14/26 (Chance of choosing a Boy [First Time]) 13/25 (Chance of choosing a Boy [Second Time]) You remove on from numerator since there is one less boy (second time), and one less student (remove one from denominator). 14/26 * 13/25 = 28% of choosing two boys.

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!